Circular No. 7364

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMETS C/2000 C2, C/2000 C3, C/2000 C4, C/2000 C5, C/2000 C6

D. A. Biesecker, SM&A Corporation and Goddard Space Flight Center, reports measurements of five comets observed with the coronagraphs aboard SOHO; the discovery observations are given below, while the astrometry and orbital elements appear on MPEC 2000-C50 through 2000-C54. Only C/2000 C6 appears to be a Kreutz sungrazer; it was first noticed by T. Lovejoy on SOHO web images, and Biesecker notes that its brightness ranged from V = 8.7 on Feb. 9.43 to 7.7 on Feb. 9.68 UT, and the comet showed a tail at 13 R_{⊙} on C3 images. The other comets showed no tail. C/2000 C2 first noted by K. Cernis, remained relatively stable in brightness (V = 6.5–6.9) during Feb. 3.70–3.84. C/2000 C3, found by Biesecker, brightened from V = 6.7 on Feb. 4.59 to 5.9 on Feb. 4.79, before fading to V = 7.0 on Feb. 5.09. C/2000 C4, found by M. Meyer, was on a trajectory closely following that of C/2000 C3, and it was assumed that the orbits are identical with a difference ΔT = 0.60 day. C/2000 C4 brightened from V = 5.9 on Feb. 5.17 to 4.9 on Feb. 5.30, before fading to V = 6.7 on Feb. 5.67. C/2000 C5, found by M. Oates, was at V = 7.5–8.0 on Feb. 7. Comets C/2000 C2 and C/2000 C5 may also be related to each other.

2000 UT α_{2000} δ_{2000} Comet
Feb. 3.704 21^{h}12.^{m}3 −15^{o}11′  C/2000 C2
4.563 21 16.4 −15 27  C/2000 C3
5.163 21 18.6 −15 15  C/2000 C4
7.788 21 28.1 −13 47  C/2000 C5
9.221 22 00.3 −17 13  C/2000 C6

AX J1740.2–2848 M. Sakano and K. Koyama, Kyoto University, write: “A post-facto

analysis of the ASCA Galactic Center Survey data obtained on 1998 Sept. 7–8 shows a coherent pulsation of 729 ± 14 s from AX J1740.2–2848 (located at α = 17^{h}40^{m}11^{s}.6, δ = −28^{o}47′48′′, equinox 2000.0; 90-percent-confidence error radius 30′′). The x-ray spectrum is fitted by a power law with photon index 0.7_{−0.5}^{+0.7}, absorbed by log N_{H} = 22.5_{−0.4}^{+0.3}. The pulsation, flat spectrum, and large absorption support AX J1740.2–2848 as a new x-ray binary pulsar near the Galactic center. The x-ray luminosity in the band 2–10 keV was 3 × 10^{34} erg/s at a distance of 8500 pc.”

2000 February 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green